a=b+3(4-c) b=5 c=2 what is a ,i need to know how the process is done to get the value of a
another question i have is how to solve y=3(x+5)(x-2) i dont remeber the steps involved in solving these equations if you could assist me it would be greatly appreciated

Answer by jim_thompson5910(35256)

a=b+3(4-c)

a=5+3(4-2)

a=5+3(2)

a=5+6

a=11

So the answer is a = 11
-------------------------------------

I'm assuming you want to find the zeros of y=3(x+5)(x-2)

y=3(x+5)(x-2)

0=3(x+5)(x-2)

3(x+5)(x-2)=0

x+5=0 or x-2=0

x = -5 or x = 2


So the solutions are x = -5 or x = 2